### v2

@@ 4.8 Verification metrics
We evaluate assimilate and forecast skill for a given variable, $\mathbf{x}_{v}$, using a latitude-weighted Root Mean Square Error (RMSE) [rasp2020weatherbench] given by
@@
Acknowledgements The authors extend their gratitude to the ECMWF and NCEP for their significant efforts to store and provide invaluable data, which are crucial for this work and the research community. Additionally, this work was carried out at the National Supercomputer Center in Tianjin, and the calculations were performed on Tianhe new generation supercomputer. We would also like to express our appreciation to the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure. This work was supported by the Science and Technology Innovation Program of Hunan Province (2022RC3070) and the National Natural Science Foundation of China (Grant Nos.  42205161, 42405146, and 42275170).
@@
Availability of data and materials All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Information. The DABench dataset is available at the Baidu Drive. The satellite observation MHS, ASCAT, AMSU-A, and SATWND observations can be accessed here.
@@
Author contribution K.J.R., L.B., B.H.D., H.Z.L., W.M.Z., J.Q.S., and W.X.W. designed the project. K.J.R., L.B., W.C.N., and B.H.D. managed and oversaw the project. W.X.W., S.M., and Y.L.Z. selected and preprocessed the observations. W.X.W., T.K.Y., and T.H. performed the model training and evaluation. W.X.W., T.H., and L.L.H. improved the model design. W.X.W., T.K.Y., X.Y.L., K.F.D., and T.H. analyze the experiment results. W.X.W., B.F., and W.C.N. wrote and revised the manuscript.

### v3

@@ 4.7 Verification metrics
We evaluate assimilation and forecast skills for a given variable, $\mathbf{x}_{v}$, using a latitude-weighted Root Mean Square Error (RMSE) [rasp2020weatherbench] given by
@@
Acknowledgements The authors extend their gratitude to the ECMWF and NCEP for their significant efforts to store and provide invaluable data, which are crucial for this work and the research community. Additionally, this work was carried out at the National Supercomputer Center in Tianjin, and the calculations were performed on Tianhe new generation supercomputer. We would also like to express our appreciation to the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure.
@@
Availability of data and materials All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Information. The DABench dataset is available at the Baidu Drive. The satellite observation MHS, ASCAT, AMSU-A, and SATWND observations can be accessed here. The TC ground truth used in this study is obtained from the IBTrACS dataset, accessible at here, which compiles recent and historical TC records from multiple agencies and provides the most comprehensive global archive of TC information.
@@
Author contribution K.J.R., L.B., B.H.D., H.Z.L., W.M.Z., J.Q.S., and W.X.W. designed the project. K.J.R., L.B., W.C.N., and B.H.D. managed and oversaw the project. W.X.W., S.M., and Y.L.Z. selected and preprocessed the observations. W.X.W., T.K.Y., and T.H. performed the model training and evaluation. W.X.W., T.H., and L.L.H. improved the model design. W.X.W., T.K.Y., X.Y.L., K.F.D., and T.H. analyzed the experimental results. W.X.W., B.F., and W.C.N. wrote and revised the manuscript.